[figure: 9.7 cm]

SEEBECK COEFFICIENT OF CePd_{3}

G.D. Mahan and M. Bartkowiak

Department of Physics and Astronomy, University of Tennessee, Knoxville, TN, 37996, USA

A brief review is provide of thermoelectric devices and materials. Most metals have a small Seebeck coefficient and are unsuitable for thermoelectric devices. The exceptions are metals with a localized magnetic moment, such as provided by a rare-earth ion. There are many rare-earth compounds which contain either cerium or ytterbium which attain Seebeck coefficients in the range of 50 − 100 μV/K. The highest reported value is found in CePd_{3} which attains a value of 125 μV/K with doping. Here we wish to explain the reason behind this larger value. Although the difference between 100 and 125 seems small, the larger value represents a breakthrough. The value of 100 μV/K appeared to be a universal maximum, so that 125 represents an achievement. A value of only 160 μV/K would make a material useful in thermoelectric devices. The high value in CePd_{3} is explained as a result of the placement of the cerium f–levels.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : G.D. Mahan

Address for correspondence : Department of Physics and Astronomy, University of Tennessee, Knoxville, TN, 37996, USA

Email address : gmahan@utkux.utcc.utk.edu